# Direct multi-model dark-matter search with gravitational-wave interferometers using data from the first part of the fourth LIGO-Virgo-KAGRA observing run

### v1

@@ IV.1 Cross-Correlation
The effect of a large population of ultralight dark photons on the interferometers is similar to a stochastic GW background, and the long coherence length of the field, e.g. $\mathcal{O}(10^{9})$ m at $m_{\rm DM}=10^{-12}\,\mathrm{eV}$ Pierce et al. (2018); Miller et al. (2021), is much larger than the separation between H1 and L1, both of which motivate a cross-correlation search. Additionally, the velocity spread of the dark photon field leads to a very narrow relative frequency spread of the order $\sim 10^{-6}$ for the signal, leading naturally to a frequency-domain peak hunt in each frequency bin of size $1/1800$ Hz (see Guo et al. (2019); Abbott et al. (2022) for previous searches with O1 and O3 data). The signal-to-noise ratio (SNR) for the frequency bin $j$ is defined as
@@
The length $T_{\text{FFT}}$ of each time segment is $1800$s, and an efficiency factor obtained from simulations is used to correct for the power loss due to this binning choice and random bin boundary when the calculated SNR is converted to upper limits Guo et al. (2019). In the presence of the dark photons, the SNR thus defined takes a negative and real value due to the relative orientation of the two detectors which results in a negative overlap reduction function, close to 0.9-0.9 for the entire frequency range we consider here, and is used in converting the SNR at each frequency bin to an upper limit. In the search for the candidate, bins that are contaminated by narrow spectral artifacts, defined to be within $\sim 0.056$ Hz, are excluded in the analysis. The upper limits are derived with the Feldman-Cousins formalism Feldman and Cousins (1998) in the absence of a detection.
@@ IV.2 BSD excess power method
In each 1-Hz band, BSD-excess-power changes the fast Fourier transform coherence time $T_{\text{FFT}}$ to match the DM signal coherence time $T_{\text{coh}}$ ensuring that all power is confined to one frequency bin for the full observing run. We then take Fourier transforms of the strain data with durations $T_{\text{FFT}}$ to make time-frequency “peakmaps” Astone et al. (2005, 2014). “Peakmaps” are collections of ones that represent when the power in particular frequency bins has exceeded a threshold in the equalized spectrum and is also a local maxima. We then count the number of peaks at each frequency to find bins with large numbers of peaks. By summing the ones in the peakmap, and not the actual power in each frequency bin, we are more robust against non-Gaussian noise artifacts that could contaminate particular frequency bins and bias our detection statistic.
We define the critical ratio $CR$ to be our detection statistic:
@@
where $y$ is the number of peaks at a particular frequency, and $\mu$ and $\sigma$ are the mean and standard deviations of the number counts across all frequencies in the band. In Gaussian noise, the $CR$ is approximately normally distributed with zero mean and unit variance, and approximately a normalized non-central $\chi^{2}$ distribution with two degrees of freedom in the presence of a signal.
We select enough candidates [^5] The number of candidates selected ranges from $20$ and $285$ at 2000 Hz and 10 Hz, respectively and depends on $T_{\text{FFT}}$. in each 1-Hz band such that one coincident candidate between two detectors would occur in Gaussian noise. This ensures that we select uniformly over the frequency range.
@@ IV.3 LPSD
The crux of the LPSD approach is to exactly match the integration time in Fourier transforms to the DM coherence time in every bin. This allows it to maximize the SNR to a theoretical maximum. The resulting necessary logarithmic frequency spacing in this method poses a challenge as FFT-based algorithms are no longer applicable. With a computational complexity scaling as $\mathcal{O}(N^{2})$, where $N$ is the number of data points, this appears to lead to intractable costs. In Göttel and Raymond (2025), a method was developed that leverages symmetries between the time and frequency domains. This allows us to perform the full logarithmic calculation with a single FFT followed by a heavily zero-suppressed transformation without the need for an expensive pre-computation.
Having efficiently calculated PSDs over large segments of data with this method, we follow the approach in Göttel et al. (2024); Göttel and Raymond (2025) where, similar to LIGO calibration, a background model is constructed by fitting cubic splines in log-log space to the PSD as a function of frequency. Because the PSDs vary by several orders of magnitude, we conduct fits over intervals of $10^{4}$ frequency bins. Within these intervals, potential DM signals are identified as positive deviations from the fits by using a profile likelihood test (see Eq. 15 in Göttel and Raymond (2025)) assuming, as validated empirically, that the residuals follow a skew-normal distribution.
@@ V.2 BSD Excess power method
Our search returned 6847 candidates present in the same frequency bin in both L1 and H1 and that did not overlap with known noise disturbances Soni et al. (2025). However, after requiring that the average $CR$ be greater than five (corresponding to $5\sigma$ significance in Gaussian noise), only ten candidates remained. To determine whether these candidates resulted from DM, we performed three follow-up procedures: (1) we varied the observation time $T_{\text{obs}}$ at the chosen $T_{\text{FFT}}$ to see if there was a steady increase in the $CR\propto T_{\text{obs}}^{1/2}$, (2) we increased $T_{\text{FFT}}>T_{\text{coh}}$ to see if there was a decrease in the $CR$ Miller et al. (2021); Vermeulen et al. (2021), and (3) we correlated the spectra from L1 and H1 with an expected DM waveform using the Wiener filter Wiener et al. (1964) and computed the residual between them Miller et al. (2022). No candidate survived all three follow-ups.
In the absence of candidates, we set upper limits on the minimum detectable strain amplitude at 95% confidence $h^{95\%}_{\rm 0,min}$ induced by a generic DM signal using Eq. 75 of Miller (2025), which depends on the detector power spectral density, the maximum of the coincident candidates’ $CRs$, $T_{\text{FFT}}$, and $T_{\text{obs}}$. Up to an $\mathcal{O}(1)$ factor that depends on the DM polarization and the geometry of LIGO, these limits can then be mapped to any kind of DM interaction that would induce a differential strain on the detector.
@@ V.3 LPSD
We obtain candidates by performing a profile likelihood ratio test in each bin, as outlined in Göttel and Raymond (2025), and require a $5\sigma$ significance corrected for the look-elsewhere effect [^6] This effect arises because when searching over a large parameter space, noise has an increased opportunity to cause a false alarm somewhere in the space. in rejecting the no-DM hypothesis. After clustering neighboring bins, requiring significance and compatibility in both detectors, we find only one candidate which was found to originate from a non-optimal setting in the background fit.
The resulting 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, where $\Lambda_{i}$ is either $\Lambda_{\rm e}$ or $\Lambda_{\gamma}$ (see Eq. 2) when assuming the other is zero, can be seen in Fig. 1(a). As one can see, we improve on the O3 results by up to an order of magnitude over the entire frequency range. We note that the O3 curve in the plot is multiplied by a factor 2.226 compared to the results in Göttel et al. (2024) after we found a conversion error. The improvement at higher frequencies can be attributed to the better detector sensitivity in O4a. At lower frequencies, a larger arm mirror thickness difference compared to O3 leads to a roughly factor of three improvement, and the gating procedure – not used in the O3 search – leads to an improvement up to a factor of two. Injection tests were performed that confirmed the upper limits at the expected level, but found a systematic leading us to overestimate the limits by 13%. This is small compared to the 30% statistical uncertainty.
@@ V.4 Upper limits
For all three pipelines, the follow-up steps, predetermined before analyzing the data, yielded no surviving candidates. Thus, we show in Fig. 1 upper limits for each of the three DM models using each pipeline, as argued in Rella (2021); Miller (2025). For all three DM models, we see significant improvements in the constraints on the coupling constants with respect to existing experiments. For dark photons, cross-correlation obtains less stringent constraints than those from BSD-excess-power and LPSD because it is significantly less sensitive to the finite-light travel time effect Morisaki et al. (2021); Manita et al. (2024). For dilatons, cross-correlation is consistent with or outperforms BSD-excess-power and LPSD because all three methods are equally sensitive to the finite light-travel time effect. For tensor bosons, the upper limits obtained by all three pipelines are consistent and significantly outperform fifth-force experiments by several orders of magnitude, as expected from Armaleo et al. (2021); Manita et al. (2023).
We note that the indirect fifth-force upper limits are stronger than those obtained from our search for scalar DM. In this case, the sensitivity of GW interferometers does not benefit significantly from the arm length, as shown in the recent GEO600 search Vermeulen et al. (2021), where the relevant effect was the differential strain induced by the oscillating beam-splitter. Improvements by accounting for the different thicknesses of the test masses Göttel et al. (2024), together with the better low-frequency sensitivity of LIGO compared to GEO600 Affeldt et al. (2014); Dooley et al. (2016), lead to substantial gains below 200 Hz relative to the GEO600 results, but still remain weaker than fifth-force constraints.
By contrast, the sensitivity to dark photons and tensor bosons in LIGO, Virgo and KAGRA depends primarily on the interferometer arm length, just as for GWs. In these cases, the differential strain induced by DM coupling to the test masses, or spacetime distortions from tensor fields, is amplified by the long arms, enhancing our sensitivity to these kinds of DM. In particular, for tensor bosons, a torsion balance responds as a single bulk object, analogous to early resonant-bar detectors, which are far less sensitive than interferometers Maggiore (2008). Thus, the sensitivity of fifth-force experiments is strongest for scalars and weakens for vectors and tensors, while GW interferometers are more powerful probes of vectors and tensors than scalars.
@@
Figure 1: Upper limits on the coupling strengths of ultralight scalar, vector and tensor bosons to standard-model particles in LIGO. Solid lines denote the constraints from this work (“cross-correlation,” “BSD excess power,” and “LPSD”). The dashed and dotted-dashed lines indicate constraints from other experiments (“Eöt-Wash” and “MICROSCOPE”) and from other searches on GW interferometer data (“GEO600”, “O3 LPSD” and “O3 acceleration effect”). (a) Our constraints apply to the coupling of dilatons to both photons and to electrons. They improve upon those from previous searches with GW interferometers Vermeulen et al. (2021); Fukusumi et al. (2023); Göttel et al. (2024) and beat those from O3 by one order of magnitude. Constraints from Eöt-Wash and MICROSCOPE correspond to the coupling of the dilaton to the electron. (b) The constraints on vector dark photon DM improve upon existing limits by a couple of order of magnitudes in the dark-photon/baryon coupling constant. To produce limits on dark photon/baryon minus lepton coupling, $U(1)_{\rm B-L}$, these limits should be multiplied by two. (c) For the first time, we derive constraints on tensor DM using GW interferometers, which are up to five orders of magnitude superior to those obtained indirectly from fifth-force experiments Sereno and Jetzer (2006); Murata and Tanaka (2015). Note that for tensor bosons, using the calculated strain amplitudes in Refs. Marzola et al. (2018); Armaleo et al. (2021) would require multiplying the limits on the tensor boson coupling constant by two.
@@ Acknowledgments
https://dcc.ligo.org/LIGO-M2300033/public. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising. We request that citations to this article use ’A. G. Abac et al. (LIGO-Virgo-KAGRA Collaboration), …’ or similar phrasing, depending on journal convention.
Appendix A Transfer functions
In the dilaton search, we use transfer functions to relate the DM-induced fluctuations of the optics and the theorized coupling constants $\Lambda_{i}^{-1}$ (see Eq. 1). Following the approach in Göttel et al. (2024), we use detailed optical simulations in Finesse to calculate the transfer functions between the interferometer’s photodiode output amplitude and DM-induced fluctuations in the beamsplitter and arm test mirrors. The results, see Fig. 2, clearly show that in O4a, LLO is the more sensitive detector, with roughly a factor three gain in sensitivity (see Section V.3) between O3 and O4a from the transfer function difference alone. At the same time, we also find that the performance of LHO has diminished over the same time period, though only slightly. In both cases, the differences are due to small changes in mirror thicknesses between the two runs.

### v3

@@ IV.1 Cross-Correlation
The effect of a large population of ultralight dark photons on the interferometers is similar to a stochastic GW background, and the long coherence length of the field, e.g. $\mathcal{O}(10^{9})$ m at $m_{\rm DM}=10^{-12}\,\mathrm{eV}$ Pierce:2018xmy; Miller:2020vsl, is much larger than the separation between H1 and L1, both of which motivate a cross-correlation search. Additionally, the velocity spread of the dark photon field leads to a very narrow relative frequency spread of the order $\sim 10^{-6}$ for the signal, leading naturally to a frequency-domain peak hunt in each frequency bin of size $1/1800$ Hz (see Guo:2019ker; LIGOScientific:2021odm for previous searches with O1 and O3 data). The signal-to-noise ratio (SNR) for the frequency bin $j$ is defined as
@@
The length $T_{\text{FFT}}$ of each time segment is $1800$s, and an efficiency factor obtained from simulations is used to correct for the power loss due to this binning choice and random bin boundary when the calculated SNR is converted to upper limits Guo:2019ker. In the presence of the dark photons, the SNR thus defined takes a negative and real value due to the relative orientation of the two detectors which results in a negative overlap reduction function, close to 0.9-0.9 for the entire frequency range we consider here, and is used in converting the SNR at each frequency bin to an upper limit. In the search for the candidate, bins that are contaminated by narrow spectral artifacts, defined to be within $\sim 0.056$ Hz, are excluded in the analysis. The upper limits are derived with the Feldman-Cousins formalism Feldman:1997qc in the absence of a detection.
@@ IV.2 BSD excess power method
In each 1-Hz band, BSD-excess-power changes the fast Fourier transform coherence time $T_{\text{FFT}}$ to match the DM signal coherence time $T_{\text{coh}}$ ensuring that all power is confined to one frequency bin for the full observing run. We then take Fourier transforms of the strain data with durations $T_{\text{FFT}}$ to make time-frequency “peakmaps” Astone:2005fj; Astone:2014esa. “Peakmaps” are collections of ones that represent when the power in particular frequency bins has exceeded a threshold in the equalized spectrum and is also a local maxima. We then count the number of peaks at each frequency to find bins with large numbers of peaks. By summing the ones in the peakmap, and not the actual power in each frequency bin, we are more robust against non-Gaussian noise artifacts that could contaminate particular frequency bins and bias our detection statistic.
We define the critical ratio CRCR to be our detection statistic:
@@
where $y$ is the number of peaks at a particular frequency, and $\mu$ and $\sigma$ are the mean and standard deviations of the number counts across all frequencies in the band. In Gaussian noise, the CRCR is approximately normally distributed with zero mean and unit variance, and approximately a normalized non-central $\chi^{2}$ distribution with two degrees of freedom in the presence of a signal.
We select enough candidates [^5] The number of candidates selected ranges from 2020 and $285$ at 2000 Hz and 10 Hz, respectively and depends on $T_{\text{FFT}}$. in each 1-Hz band such that one coincident candidate between two detectors would occur in Gaussian noise. This ensures that we select uniformly over the frequency range.
@@ IV.3 LPSD
The crux of the LPSD approach is to exactly match the integration time in Fourier transforms to the DM coherence time in every bin. This allows it to maximize the SNR to a theoretical maximum. The resulting necessary logarithmic frequency spacing in this method poses a challenge as FFT-based algorithms are no longer applicable. With a computational complexity scaling as $\mathcal{O}(N^{2})$, where NN is the number of data points, this appears to lead to intractable costs. In Gottel:2025aqm, a method was developed that leverages symmetries between the time and frequency domains. This allows us to perform the full logarithmic calculation with a single FFT followed by a heavily zero-suppressed transformation without the need for an expensive pre-computation.
Having efficiently calculated PSDs over large segments of data with this method, we follow the approach in Gottel:2024cfj; Gottel:2025aqm where, similar to LIGO calibration, a background model is constructed by fitting cubic splines in log-log space to the PSD as a function of frequency. Because the PSDs vary by several orders of magnitude, we conduct fits over intervals of $10^{4}$ frequency bins. Within these intervals, potential DM signals are identified as positive deviations from the fits by using a profile likelihood test (see Eq. 15 in Gottel:2025aqm) assuming, as validated empirically, that the residuals follow a skew-normal distribution.
@@ V.2 BSD Excess power method
Our search returned 6847 candidates present in the same frequency bin in both L1 and H1 and that did not overlap with known noise disturbances LIGO:2024kkz. However, after requiring that the average CRCR be greater than five (corresponding to $5\sigma$ significance in Gaussian noise), only ten candidates remained. To determine whether these candidates resulted from DM, we performed three follow-up procedures: (1) we varied the observation time $T_{\text{obs}}$ at the chosen $T_{\text{FFT}}$ to see if there was a steady increase in the $CR\propto T_{\text{obs}}^{1/2}$, (2) we increased $T_{\text{FFT}}>T_{\text{coh}}$ to see if there was a decrease in the CRCR Miller:2020vsl; Vermeulen:2021epa, and (3) we correlated the spectra from L1 and H1 with an expected DM waveform using the Wiener filter wiener1964extrapolation and computed the residual between them Miller:2022wxu. No candidate survived all three follow-ups.
In the absence of candidates, we set upper limits on the minimum detectable strain amplitude at 95% confidence $h^{95\%}_{\rm 0,min}$ induced by a generic DM signal using Eq. 75 of Miller:2025yyx, which depends on the detector power spectral density, the maximum of the coincident candidates’ CRsCRs, $T_{\text{FFT}}$, and $T_{\text{obs}}$. Up to an $\mathcal{O}(1)$ factor that depends on the DM polarization and the geometry of LIGO, these limits can then be mapped to any kind of DM interaction that would induce a differential strain on the detector.
@@ V.3 LPSD
We obtain candidates by performing a profile likelihood ratio test in each bin, as outlined in Gottel:2025aqm, and require a $5\sigma$ significance corrected for the look-elsewhere effect [^6] This effect arises because when searching over a large parameter space, noise has an increased opportunity to cause a false alarm somewhere in the space. in rejecting the no-DM hypothesis. After clustering neighboring bins, requiring significance and compatibility in both detectors, we find only one candidate which was found to originate from a non-optimal setting in the background fit.
The resulting 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, where $\Lambda_{i}$ is either $\Lambda_{\rm e}$ or $\Lambda_{\gamma}$ (see Eq. 2) when assuming the other is zero, can be seen in Fig. 1(a). As one can see, we improve on the O3 results by up to an order of magnitude over the entire frequency range. We note that the O3 curve in the plot is multiplied by a factor 2.226 compared to the results in Gottel:2024cfj after we found a conversion error. The improvement at higher frequencies can be attributed to the better detector sensitivity in O4a. At lower frequencies, a larger arm mirror thickness difference compared to O3 leads to a roughly factor of three improvement, and the gating procedure – not used in the O3 search – leads to an improvement up to a factor of two. Injection tests were performed that confirmed the upper limits at the expected level, but found a systematic leading us to overestimate the limits by 13%. This is small compared to the 30% statistical uncertainty.
@@ V.4 Upper limits
Figure 1: Upper limits on the coupling strengths of ultralight scalar, vector and tensor bosons to standard-model particles in LIGO. Solid lines denote the constraints from this work (“cross-correlation,” “BSD excess power,” and “LPSD”). The dashed and dotted-dashed lines indicate constraints from other experiments (“Eöt-Wash” and “MICROSCOPE”) and from other searches on GW interferometer data (“GEO600”, “O3 LPSD” and “O3 acceleration effect”). (a) Our constraints apply to the coupling of dilatons to both photons and to electrons. They improve upon those from previous searches with GW interferometers Vermeulen:2021epa; Fukusumi:2023kqd; Gottel:2024cfj and beat those from O3 by one order of magnitude. Constraints from Eöt-Wash Su:1994gu; Schlamminger:2007ht and MICROSCOPE Touboul:2012ui; Berge:2017ovy correspond to the coupling of the dilaton to the electron. (b) The constraints on vector dark photon DM improve upon existing limits by a couple of order of magnitudes in the dark-photon/baryon coupling constant. To produce limits on dark photon/baryon minus lepton coupling, $U(1)_{\rm B-L}$, these limits should be multiplied by two. (c) For the first time, we derive constraints on tensor DM using GW interferometers, which are up to five orders of magnitude superior to those obtained indirectly from fifth-force experiments Sereno:2006mw; Murata:2014nra. Note that for tensor bosons, using the calculated strain amplitudes in Refs. Marzola:2017lbt; Armaleo:2020efr would require multiplying the limits on the tensor boson coupling constant by two.
For all three pipelines, the follow-up steps, predetermined before analyzing the data, yielded no surviving candidates. Thus, we show in Fig. 1 upper limits for each of the three DM models using each pipeline, as argued in rellathesis; Miller:2025yyx. For all three DM models, we see significant improvements in the constraints on the coupling constants with respect to existing experiments. For dark photons, cross-correlation obtains less stringent constraints than those from BSD-excess-power and LPSD because it is significantly less sensitive to the finite-light travel time effect Morisaki:2020gui; Manita:2023mnc. For dilatons, cross-correlation is consistent with or outperforms BSD-excess-power and LPSD because all three methods are equally sensitive to the finite light-travel time effect. For tensor bosons, the upper limits obtained by all three pipelines are consistent and significantly outperform fifth-force experiments by several orders of magnitude, as expected from Armaleo:2020efr; Manita:2022tkl.
We note that the indirect fifth-force upper limits are stronger than those obtained from our search for scalar DM. In this case, the sensitivity of GW interferometers does not benefit significantly from the arm length, as shown in the recent GEO600 search Vermeulen:2021epa, where the relevant effect was the differential strain induced by the oscillating beam-splitter. Improvements by accounting for the different thicknesses of the test masses Gottel:2024cfj, together with the better low-frequency sensitivity of LIGO compared to GEO600 Affeldt:2014rza; Dooley:2015fpa, lead to substantial gains below 200 Hz relative to the GEO600 results, but still remain weaker than fifth-force constraints.
By contrast, the sensitivity to dark photons and tensor bosons in LIGO, Virgo and KAGRA depends primarily on the interferometer arm length, just as for GWs. In these cases, the differential strain induced by DM coupling to the test masses, or spacetime distortions from tensor fields, is amplified by the long arms, enhancing our sensitivity to these kinds of DM. In particular, for tensor bosons, a torsion balance responds as a single bulk object, analogous to early resonant-bar detectors, which are far less sensitive than interferometers maggiore2008gravitational. Thus, the sensitivity of fifth-force experiments is strongest for scalars and weakens for vectors and tensors, while GW interferometers are more powerful probes of vectors and tensors than scalars.
@@ VI Conclusions
Appendix A Transfer functions
In the dilaton search, we use transfer functions to relate the DM-induced fluctuations of the optics and the theorized coupling constants $\Lambda_{i}^{-1}$ (see Eq. 1). Following the approach in Gottel:2024cfj, we use detailed optical simulations in Finesse to calculate the transfer functions between the interferometer’s photodiode output amplitude and DM-induced fluctuations in the beamsplitter and arm test mirrors. The results, see Fig. 2, clearly show that in O4a, LLO is the more sensitive detector, with roughly a factor three gain in sensitivity (see Section V.3) between O3 and O4a from the transfer function difference alone. At the same time, we also find that the performance of LHO has diminished over the same time period, though only slightly. In both cases, the differences are due to small changes in mirror thicknesses between the two runs.
@@ Acknowledgments
J.S. is supported by the Peking University under startup Grant No. 7101302974, the NSFC under Grants No. 12025507, No. 12150015, No.12450006.
@@
https://dcc.ligo.org/LIGO-M2300033/public. For the purpose of open access, the authors have applied a Creative Commons Attribution ($C$ BY) license to any Author Accepted Manuscript version arising. We request that citations to this article use ’A. G. Abac et al. (LIGO-Virgo-KAGRA Collaboration), …’ or similar phrasing, depending on journal convention.
The LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA Collaboration
The LIGO Scientific Collaboration, the Virgo Collaboration, and the KAGRA Collaboration